OpenStudy (dkdkdkdkdkdkdkdkdk91919288282x):

The amount of an ordinary 7500 annuity for 3 years at 12 percent compounded quarterly is?

OpenStudy (maketheworldyourstudygroup):

$7500.00[1+{(12/100)/4}]4*3= $10,693.21 as the formula stands like A=P(1+rn)nt where A = Amount. P= Principal r=rate of interest expressed as percentage n=number of times the interest is compounded in a year t=total time